# 2cm2.5cm2.5cm2cm

### v2

@@ 3.2 Monte Carlo Simulation
Table 1 summarises the results obtained for settings I and II. Empirical means and standard deviation (SD) of the CMLEs are provided by sample size. TResults show adequate performance of the estimation method, with empirical Monte Carlo means close to the values and standard deviation decreasing with the increase in sample size, as expected.
@@
Figure 1 illustrates the asymptotic normality of the CMLEs for setting I. Non-parametric density estimator plots of the standardized parameter estimates are displayed alongside the standard Gaussian density (solid line). Dashed and dotted curves indicate densities estimated from the experiments carried out with $n=500$ and $n=1000$, respectively. The density curves of the parameter estimates are mostly overlapping, but some improvement can be seen with the increase of the sample size. Additional density plots for the other parameter configurations can be found in the Supplementary Material.
Figure 1: Non-parametric density plots of standardized parameter estimates due to Setting I under sample sizes $n=500$ and $n=1000$. The solid line corresponds to the standard Gaussian density function.
@@ 4 Testing Constant Dispersion
Testing constant value of the dispersion parameter, i.e. $\phi_{t}=\alpha_{0}$ for all $t$ under model (3), is equivalent to testing $H_{0}:\alpha_{1}=\alpha_{2}=0$ against $H_{1}:\alpha_{1}\neq 0\,\,\mbox{or}\,\,\alpha_{2}\neq 0$. Note that the null hypothesis belongs to the boundary of the parameter space, which is a non-standard testing problem; for instance, see Self and Liang (1987), Andrews (2001), and Crainiceanu and Ruppert (2004). We develop and compare heuristically two parametric bootstrap methods; the classical or unrestricted, and the restricted bootstrap recently developed by Cavaliere et al. (2016). The first method considers the usual parametric bootstrap (Efron and Tibshirani, 1993) replications based on the unrestricted CMLEs, while the latter uses the CMLEs under $H_{0}$. Algorithm 1 describes the estimation of the test’s p-value with $B$ replications of the restricted or unrestricted bootstrap.
@@
      3A) $\boldsymbol{Y}_{b}\sim\mbox{tv NB-INGARCH}(\widehat{\boldsymbol{\theta}}_{\mathcal{H}_{0}})$;
@@
      3B) $\boldsymbol{Y}_{b}\sim\mbox{tv NB-INGARCH}(\widehat{\boldsymbol{\theta}})$;
@@
      4. Obtain $\widehat{\boldsymbol{\theta}}_{\mathcal{H}_{0}}^{b}$ and $\boldsymbol{\widehat{\theta}}^{b}$ fitting tv NB-INGARCH models to $\boldsymbol{Y}_{b}$ under the null and alternative hypothesis;
@@ 6. If $p_{B}=\sum_{b=1}^{B}I\{LR^{b}>LR\}/B<\alpha$ reject $\mathcal{H}_{0}$;
Algorithm 1 Bootstrap likelihood ratio test of constant ($\mathcal{H}_{0}:\alpha_{1}=\alpha_{2}=0$) versus time-varying dispersion ($\mathcal{H}_{1}:\alpha_{1}\neq 0\mbox{ or }\alpha_{2}\neq 0$) for a tv NB-INGARCH model. Alternatives to step 3 yield restricted (3A) or unrestricted bootstrap (3B) estimators of the test’s p-value, $p_{B}$, where $B$ is the number of replications.
We use a Monte Carlo simulation study to investigate how these methods achieve the desirable significance levels. Time series from the NB tv-DINGARCH process with 200 time observations are simulated under the null hypothesis using four different (varying) settings of the parameter vector $(\beta_{0},\beta_{1},\beta_{2},\alpha_{0})^{\top}$ as follows: (C1) $\beta_{0}=2,\alpha_{0}=1$, $\beta_{1}=0.4,\beta_{2}=0.3$, (C2) $\beta_{0}=2,\alpha_{0}=1$, $\beta_{1}=0,\beta_{2}=0$, (C3) $\beta_{0}=3,\alpha_{0}=0.5$, $\beta_{1}=0.3,\beta_{2}=0.4$, and (C4) $\beta_{0}=3,\alpha_{0}=0.5$, $\beta_{1}=0,\beta_{2}=0$. For each configuration, 500 Monte Carlo replications are used to calculate the empirical significance levels by employing the competing methodologies. The number of replications used to estimate bootstrap $p$-values is $B=500$. Table 2 displays the proportion of times that the restricted and unrestricted bootstrap procedures rejected the null hypothesis. Parameter configurations are set in a way that C2 and C4 are variations of C1 and C3 that do not include effects on the mean. Additional evidence is provided in the Supplementary Material.
@@
Table 2: Nominal significance levels produced by the restricted and unrestricted bootstrap hypothesis tests for $H_{0}:\alpha_{1}=\alpha_{2}=0$ for various parameter configurations.
Notably, the restricted parametric bootstrap agrees with the set significance levels, something that occurs under the four parameter configurations investigated in this study, while the unrestricted parametric bootstrap does not provide satisfactory results by underestimating the significance levels. These results show the importance of considering a restricted bootstrap for testing constant dispersion in the tv-DINGARCH models.
@@ 5 Measles Data Analysis
The linear negative binomial tv-DINGARCH model is now applied to modeling the weekly number of reported measles infections in North Rhine-Westphalia, Germany. The series is observed between January 2001 and May 2013 (646 observations), publicly available in the R package tscount. Figure 2 displays $\{Y_{t}\}_{t=1}^{646}$ on the left, and the series autocorrelation function on the right.
@@
Fitting of the NB tv-DINGARCH model will be compared to the fit of ordinary NBINGARCH model with an identical mean time series structure. The purpose of this exercise is to assess how a time-varying dispersion changes the model adequacy to this real data. To this end, we will evaluate goodness-of-fit and predictions for each model. $\mathcal{M}_{tv}$ will denote the NB tv-DINGARCH model while $\mathcal{M}_{ord}$ denotes the ordinary case (which assumes a constant dispersion).
@@
Table 3: Conditional maximum likelihood estimates, standard errors (se), and 95% confidence intervals for log-linear tv-DINGARCH and ordinary log-linear INGARCH models applied to weekly counts of measles in Germany.
In addition, we consider Probability Integral Transform (PIT) plots (Czado et al., 2009), an approach that enables the comparison of count data models via their predictive distributions. A model providing a good fit to the data in this aspect will render a PIT plot resembling a uniform distribution, where major deviations typically indicate problems of overdispersion or underdispersion of the model’s predictive distribution. These are reported in Figure 3 for both models, and while that of $\mathcal{M}_{tv}$ is near the uniform as desired, the upside-down U shaped PIT of $\mathcal{M}_{ord}$ points to a non adequate fit.
Figure 3: Probability Integral Transform (PIT) assessment for $\mathcal{M}_{ord}$ (on the left) and $\mathcal{M}_{tv}$ (to the right).
It is often of interest to practitioners to select between models according to their forecasting power. To that end, we consider a forecasting exercise that explroes one-step-ahead (OSA) prediction from $\mathcal{M}_{tv}$ and $\mathcal{M}_{ord}$. This is done in a recursive manner in order to provide multiple OSA predictions from each model. We start by defining the initial training data from the beginning of the study until October 2004 (week 4) which contains 200 observations. Both models are fitted to the training set and used to predict the next week’s counts (week 1 of November 2004) via the conditional median and mode of the distributions. The conditional (or predictive) distributions are Negative-Binomial($\hat{\mu}_{t+1},\hat{\phi}_{t+1}$), or Negative-Binomial($\hat{\mu}_{t+1},\hat{\phi}$) respectively for $\mathcal{M}_{tv}$ and $\mathcal{M}_{ord}$ and $t=200$ at this step. Once the prediction is obtained, we add week 1 of November 2004 to the training set, refit both models and gather the new OSA predictions. Proceeding until the end of the study period gives a total of 446 predictions from each model. Pseudocode describing the steps to this prediction exercise is given in Algorithm 2.
@@ 1. $\boldsymbol{Y}(s)\leftarrow Y_{[1:s]}$;
      2. Fit the tv-NBINGARCH model to $\boldsymbol{Y}(s)$;
      3. From 2, gather the CMLEs $\widehat{\boldsymbol{\theta}}(s)$ and the fitted $\widehat{\boldsymbol{\lambda}}(s)$ and $\widehat{\boldsymbol{\phi}}(s)$ of step $s$;
@@ 2. Fit the tv-NBINGARCH model to $\boldsymbol{Y}(s)$;
      5A) $\widehat{Y}_{s+1}=\widehat{\mu}_{s+1}$
       // prediction via mean
      5B) $\widehat{Y}_{s+1}=\texttt{qbinom}(0.5,\texttt{size}=\widehat{\phi}_{s+1},\texttt{mu}=\widehat{\mu}_{s+1})$
@@
Algorithm 2 Recursive algorithm for obtaining $\boldsymbol{\widehat{Y}}$, the one-step-ahead (OSA) predicted values of $\boldsymbol{Y}_{[n_{0}+1:n]}$. The training data at iteration $s$, $\boldsymbol{Y}(s)$, is incremented and the model is refitted to obtain the OSA forecast $\widehat{Y}_{s+1}$. Steps 5A) and 5B) provide alternative prediction methods via the mean or median.
Algorithm 2 is presented in terms of the $\mathcal{M}_{tv}$ model but works similarly for $\mathcal{M}_{ord}$. In this case, $\mathcal{M}_{ord}$ is fitted in step two and the predictions in 5A) and 5B) take in the (fixed) dispersion parameter estimated in step 2. By employing the Algorithm with $\mathcal{M}_{tv}$ and $\mathcal{M}_{ord}$, their predictions are summarized via the root mean forecasting error (RMSFE). Let $n_{0}$ denote the time point chosen to start the prediction exercise; in our case $n_{0}=200$ the RMSFE of the forecasting step $t$ is
@@
where $Y_{s}$ and $\widehat{Y}_{s}$, are the observed and predicted counts, respectively. Calculation of RMSFE, from the 446 total predictions from $\mathcal{M}_{tv}$ and $\mathcal{M}_{ord}$, yields Figure 4. On the left, $\widehat{Y}_{s}$ is the mode of the conditional distribution and on the right it is taken as the median. Prediction by the median yields a smaller prediction error in comparison to the mode, and in both cases the tv-DINGARCH model produces the smallest RMSFE values.
Figure 4: RMSFE of predictions obtained with the fit of ordinary (dashed lines) and time-varying dispersion (solid lines) INGARCH models to the weekly count of measles in North Rhine-Westphalia, Germany. On the left, predicted values are the mode of the predictive distributions, whereas the median is taken as the point on the right.
This empirical illustration on the weekly number of measles cases in North Rhine-Westphalia, Germany, demonstrated that the time-varying dispersion is a promising and important extension of the ordinary INGARCH processes that achieved improvement in terms of goodness-of-fit and forecasting.
@@ 6 Conclusions
We proposed a class of time-varying dispersion INGARCH (tv-DINGARCH) models and explore stochastic properties such as stationarity and ergodicity. Estimation of parameters was addressed through conditional maximum likelihood estimation (CMLE) and its associated asymptotic theory was established. Monte Carlo simulations were conducted to evaluate the performance of the CMLE. Moreover, we developed bootstrap methodologies to test for constant dispersion and showed via simulated studies that the restricted bootstrap is preferred over the unrestricted parametric one. We analyzed the weekly number of reported measles cases in north Rhine-Westphalia, Germany, from January 2001 to May 2013, and found that the tv-DINGARCH approach delivers much better results regarding goodness-of-fit and prediction when compared to the ordinary INGARCH model.
@@
Let $w$ be a continuous function with compact support on $(0,\infty)^{2}$ and assumed to be bounded $|w|<1$ without loss of generality. Consider $k=1$, ${\bf x},{\bf z}\in(0,\infty)^{2}$, and $\epsilon>0$ arbitrary. Denote by $f^{mp}_{\boldsymbol{\xi}}(\cdot)$ the probability function of a mixed Poisson distribution with mean $\lambda$ and variance $\lambda+\lambda^{2}/\phi$, where $\boldsymbol{\xi}=(\lambda,\phi)^{\top}$. It follows that
@@
$\sum_{y=0}^{\infty}|w(\boldsymbol{\tau}+{\bf B}(y,y)^{\top}+{\bf A}\boldsymbol{\bf x})f^{mp}_{\bf x}(y)-w(\boldsymbol{\tau}+{\bf B}(y,y)^{\top}+{\bf A}\boldsymbol{\bf z})f^{mp}_{\bf z}(y)|\leq$
$\sum_{y=0}^{\infty}f^{mp}_{\bf x}(y)|w(\boldsymbol{\tau}+{\bf B}(y,y)^{\top}+{\bf A}\boldsymbol{\bf x})-w(\boldsymbol{\tau}+{\bf B}(y,y)^{\top}+{\bf A}\boldsymbol{\bf z})|+$ (A-1)
$\sum_{y=0}^{\infty}|w(\boldsymbol{\tau}+{\bf B}(y,y)^{\top}+{\bf A}\boldsymbol{\bf z})|\,|f^{mp}_{\bf x}(y)-f^{mp}_{\bf z}(y)|.$ (A-2)
We now find an upper bound for the term (A-2). Denote by $f^{pois}_{\lambda}(\cdot)$ the probability function of a Poisson distribution with mean $\lambda$. From the mixed Poisson stochastic representation, we have that $f^{mp}_{\bf x}(y)=E(f^{pois}_{x_{1}Z_{1}}(y))$, where $Z_{1}$ is the associated latent random variables with distribution depending on $x_{2}$. Similar representation holds for $f^{mp}_{\bf z}(y)$ in terms of an associated latent factor $Z_{2}$ with distribution depending on $z_{2}$. Using this and the recalling the fact that $|g|<1$, we obtain that
$\sum_{y=0}^{\infty}|w(\boldsymbol{\tau}+{\bf B}(y,y)^{\top}+{\bf A}\boldsymbol{\bf z})|\,|f^{mp}_{\bf x}(y)-f^{mp}_{\bf z}(y)|\leq\sum_{y=0}^{\infty}|f^{mp}_{\bf x}(y)-f^{mp}_{\bf z}(y)|=$
$\sum_{y=0}^{\infty}|E(f^{pois}_{x_{1}Z_{1}}(y))-E(f^{pois}_{z_{1}Z_{2}}(y))|\leq E\left(\sum_{y=0}^{\infty}|f^{pois}_{x_{1}Z_{1}}(y)-f^{pois}_{x_{2}Z_{2}}(y)|\right).$ (A-3)
By using inequality (A.1) from Wang et al. (2014) (see also Liu (2012)), we have that $\sum_{y=0}^{\infty}|f^{pois}_{x_{1}Z_{1}}(y)-f^{pois}_{x_{2}Z_{2}}(y)|\leq 2\left(1-\exp\{-|x_{1}Z_{1}-x_{2}Z_{2}|\}\right)$. Hence, (A-1) is bounded above by
$E\left(\sum_{y=0}^{\infty}|f^{pois}_{x_{1}Z_{1}}(y)-f^{pois}_{x_{2}Z_{2}}(y)|\right)\leq 2E\left(1-e^{-|x_{1}Z_{1}-x_{2}Z_{2}|}\right)\leq$
$2E\left(1-e^{-|x_{1}-x_{2}|(Z_{1}+Z_{2})}\right)\leq 2E\left(|x_{1}-x_{2}|(Z_{1}+Z_{2})\right)=2|x_{1}-x_{2}|\leq 2\|{\bf x}-{\bf z}\|_{p},$

### v3

@@ 3.2 Monte Carlo Simulation
Table 1 summarises the results obtained for settings I and II. Empirical means and standard deviation (SD) of the CMLEs are reported by sample size. The results show adequate performance of the estimation method, with empirical Monte Carlo means close to the values and standard deviation decreasing with the increase in sample size, as expected.
@@
Figure 1 illustrates the asymptotic normality of the CMLEs for setting I. Non-parametric density estimator plots of the standardized parameter estimates are displayed alongside the standard Gaussian density (solid line). Dashed and dotted curves indicate densities estimated from the experiments carried out with $n=500$ and $n=1000$, respectively. The density curves of the parameter estimates are mostly overlapping, but some improvement can be seen with the increase in the sample size. Additional density plots for the other parameter configurations can be found in the Supplementary Material.
Figure 1: Non-parametric density plots of standardized parameter estimates due to Setting I when $n=500$ and $n=1000$. The solid line corresponds to the standard Gaussian density function.
@@ 4 Testing Constant Dispersion
Testing constant value of the dispersion parameter, i.e. $\phi_{t}=\alpha_{0}$ for all $t$ under model (3), is equivalent to testing $H_{0}:\alpha_{1}=\alpha_{2}=0$ against $H_{1}:\alpha_{1}>0\,\,\mbox{or}\,\,\alpha_{2}>0$. Note that the null hypothesis implies that the parameters $\alpha_{1}$ and $\alpha_{2}$ belong to the boundary of the parameter space. Hence, this is a non-standard testing problem; see, for instance, Self and Liang (1987), Andrews (2001), and Crainiceanu and Ruppert (2004).
We consider the likelihood ratio test to develop and compare heuristically two parametric bootstrap methods; the classical or unrestricted, and the restricted bootstrap recently developed by Cavaliere et al. (2016). The first method considers the usual parametric bootstrap (Efron and Tibshirani, 1993) replications based on the unrestricted CMLEs, while the latter uses the CMLEs under $H_{0}$. Algorithm 1 specifies computation of the likelihood ratio test’s p-value with $B$ replications of the restricted or unrestricted bootstrap.
@@
      3A) $\boldsymbol{Y}_{b}\sim\mbox{NB tv-DINGARCH}(\widehat{\boldsymbol{\theta}}_{\mathcal{H}_{0}})$;
@@
      3B) $\boldsymbol{Y}_{b}\sim\mbox{NB tv-DINGARCH}(\widehat{\boldsymbol{\theta}})$;
@@
      4. Obtain $\widehat{\boldsymbol{\theta}}_{\mathcal{H}_{0}}^{b}$ and $\boldsymbol{\widehat{\theta}}^{b}$ fitting NB tv-DINGARCH models to $\boldsymbol{Y}_{b}$ under the null and alternative hypothesis;
@@ 6. If $p_{B}=\sum_{b=1}^{B}I\{LR^{b}>LR\}/B<\alpha$ reject $\mathcal{H}_{0}$;
Algorithm 1 Bootstrap likelihood ratio test of constant ($\mathcal{H}_{0}:\alpha_{1}=\alpha_{2}=0$) versus time-varying dispersion ($\mathcal{H}_{1}:\alpha_{1}>0\mbox{ or }\alpha_{2}>0$) for a NB tv-DINGARCH model. Alternatives to step 3 yield restricted (3A) or unrestricted bootstrap (3B) estimators of the test’s p-value, $p_{B}$, where $B$ is the number of replications.
We use a Monte Carlo simulation study to investigate how these methods achieve the desirable significance levels. Time series from the NB tv-DINGARCH process with 200 observations are simulated under the null hypothesis using four different (varying) settings of the parameter vector $(\beta_{0},\beta_{1},\beta_{2},\alpha_{0})^{\top}$ as follows: (C1) $\beta_{0}=2,\alpha_{0}=1$, $\beta_{1}=0.4,\beta_{2}=0.3$, (C2) $\beta_{0}=2,\alpha_{0}=1$, $\beta_{1}=0,\beta_{2}=0$, (C3) $\beta_{0}=3,\alpha_{0}=0.5$, $\beta_{1}=0.3,\beta_{2}=0.4$, and (C4) $\beta_{0}=3,\alpha_{0}=0.5$, $\beta_{1}=0,\beta_{2}=0$. For each configuration, 500 Monte Carlo replications are used to calculate the empirical significance levels by employing the competing methodologies. The number of replications used to estimate bootstrap $p$-values is $B=500$. Table 2 displays the proportion of times that the restricted and unrestricted bootstrap procedures rejected the null hypothesis. Parameter configurations are set in a way that C2 and C4 are variations of C1 and C3 that do not include effects on the mean. Additional evidence is provided in the Supplementary Material.
@@
Table 2: Achieved significance levels of likelihood ratio test employing Algorithm 1 for various parameter configurations.
Notably, the restricted parametric bootstrap achieves levels close to nominal significance levels, something that occurs under all four parameter configurations investigated in this study. The unrestricted parametric bootstrap does not provide satisfactory results and it underestimates the significance levels. Hence, in applications, it is important to employ a restricted bootstrap for testing constant dispersion in the tv-DINGARCH models.
The results obtained from both restricted and unrestricted bootstrap were already expected. As discussed by Cavaliere et al. (2016), it is well-known that the unrestricted (classical) bootstrap does not apply when one parameter belongs to the boundary of the parameter space. This is indeed the case where the null hypothesis is that $H_{0}:\alpha_{1}=\alpha_{2}=0$, both parameters belong to the boundary space. The restricted bootstrap proposed by Cavaliere et al. (2016) overcomes this issue and provides valid results as theoretically proven in their paper and empirically illustrated here for DINGARCH models.
@@ 5 Measles Data Analysis
The linear negative binomial tv-DINGARCH model is now applied to modeling the weekly number of reported measles infections in North Rhine-Westphalia, Germany. The series is observed between January 2001 and May 2013 (646 observations). These data are publicly available in the R package tscount. Figure 2 displays $\{Y_{t}\}_{t=1}^{646}$ on the left, and the series autocorrelation function on the right.
@@
The fitting of the NB tv-DINGARCH model will be compared to the fit of the ordinary NB INGARCH model with an identical mean time series structure. The purpose of this exercise is to assess how a time-varying dispersion changes the model adequacy to this real data. To this end, we will evaluate goodness-of-fit and predictions for each model, i.e. $\mathcal{M}_{tv}$ will denote the NB tv-DINGARCH model while $\mathcal{M}_{ord}$ denotes the ordinary case (which assumes a constant dispersion).
@@
Table 3: Conditional maximum likelihood estimates, standard errors (se), and 95% confidence intervals for $\mathcal{M}_{tv}$ and $\mathcal{M}_{ord}$ models applied to weekly counts of measles in Germany.
In addition, we consider Probability Integral Transform (PIT) plots (Czado et al., 2009), an approach that enables the comparison of count time series models via their predictive distributions. A model providing a good fit to the data in this aspect will render a PIT plot resembling a uniform distribution, where major deviations typically indicate problems of overdispersion or underdispersion of the model’s predictive distribution. These graphs are shown in Figure 3 for both models and while the plot for $\mathcal{M}_{tv}$ resembles uniformity, as desired, the upside-down U-shaped PIT of $\mathcal{M}_{ord}$ indicates a non adequate fit.
Figure 3: PIT plots for $\mathcal{M}_{ord}$ (left) and $\mathcal{M}_{tv}$ (right).
It is often of interest to practitioners to select between models according to their forecasting power. To that end, we consider a forecasting exercise that explores one-step-ahead (OSA) prediction from $\mathcal{M}_{tv}$ and $\mathcal{M}_{ord}$. This is done in a recursive manner to provide multiple OSA predictions from each model. We start by defining the initial training data from the beginning of the study until October 2004 (week 4) which contains 200 observations. Both models are fitted to the training set and used to predict the next week’s counts (week 1 of November 2004) via the conditional median and mode of the distributions. The conditional (or predictive) distributions are $\mbox{NB}(\hat{\mu}_{t+1},\hat{\phi}_{t+1})$ and $\mbox{NB}(\hat{\mu}_{t+1},\hat{\phi})$ respectively for $\mathcal{M}_{tv}$ and $\mathcal{M}_{ord}$ and $t=200$ at this step. Once the prediction is obtained, we add week 1 of November 2004 to the training set, refit both models, and gather the new OSA predictions. Proceeding until the end of the study period gives a total of 446 predictions from each model. A pseudocode describing the steps to this prediction exercise is given in Algorithm 2.
@@ 1. $\boldsymbol{Y}(s)\leftarrow Y_{[1:s]}$;
      2. Fit the NB tv-DINGARCH model to $\boldsymbol{Y}(s)$;
      3. From 2, put together the CMLEs $\widehat{\boldsymbol{\theta}}(s)$ and the fitted $\widehat{\boldsymbol{\lambda}}(s)$ and $\widehat{\boldsymbol{\phi}}(s)$ of step $s$;
@@ 2. Fit the NB tv-DINGARCH model to $\boldsymbol{Y}(s)$;
      5A) Compute the mode of the $\mbox{NB}(\widehat{\mu}_{s+1},\widehat{\phi}_{s+1})$ distribution and denote it by $\widehat{Y}_{s+1}$;
       // prediction via mode
      5B) Compute the median of the $\mbox{NB}(\widehat{\mu}_{s+1},\widehat{\phi}_{s+1})$ distribution and denote it by $\widehat{Y}_{s+1}$;
@@
Algorithm 2 Recursive algorithm for obtaining $\boldsymbol{\widehat{Y}}$, the one-step-ahead (OSA) predicted values of $\boldsymbol{Y}_{[n_{0}+1:n]}$. The training data at iteration $s$, $\boldsymbol{Y}(s)$, is incremented and the model is refitted to obtain the OSA forecast $\widehat{Y}_{s+1}$. Steps 5A) and 5B) provide alternative prediction methods via the mode and median.
Algorithm 2 is presented in terms of the $\mathcal{M}_{tv}$ model but applies similarly for the case of $\mathcal{M}_{ord}$. In this case, $\mathcal{M}_{ord}$ is fitted in step two, and the predictions in 5A) and 5B) take in the (fixed) dispersion parameter estimated in step 2. By employing the Algorithm with $\mathcal{M}_{tv}$ and $\mathcal{M}_{ord}$, their predictions are summarized via the root mean square forecasting error (RMSFE). Let $n_{0}$ denote the time point chosen to start the prediction exercise; in this case $n_{0}=200$. The RMSFE of the forecasting step $t$ is
@@
where $Y_{s}$ and $\widehat{Y}_{s}$, are the observed and predicted counts, respectively. Calculation of RMSFE, from the 446 total predictions from $\mathcal{M}_{tv}$ and $\mathcal{M}_{ord}$, yields Figure 4. On the left, $\widehat{Y}_{s}$ is the mode of the conditional distribution and on the right, it is taken as the median. Prediction by the median yields a smaller prediction error in comparison to the mode, and in both cases the tv-DINGARCH model produces the smallest RMSFE values.
Figure 4: RMSFE of predictions obtained with the fit of ordinary (dashed lines) and time-varying dispersion (solid lines) INGARCH models to the weekly count of measles in North Rhine-Westphalia, Germany. On the left, predicted values are the mode of the predictive distributions, whereas the median is taken on the right.
The analysis of the weekly number of measles cases in North Rhine-Westphalia, Germany, demonstrated that the time-varying dispersion is a promising and important extension of the ordinary INGARCH processes that achieved improvement in terms of goodness-of-fit and forecasting.
@@ 6 Conclusions
We proposed a class of time-varying dispersion INGARCH (tv-DINGARCH) models and explored stochastic properties such as stationarity and ergodicity. Estimation of parameters was addressed through conditional maximum likelihood estimation (CMLE) and its associated asymptotic theory was established. Monte Carlo simulations were conducted to evaluate the performance of the CMLE. Moreover, we developed bootstrap methodologies to test for constant dispersion and showed via simulated studies that the restricted bootstrap is preferred over the unrestricted parametric one. We analyzed the weekly number of reported measles cases in North Rhine-Westphalia, Germany, from January 2001 to May 2013, and found that the tv-DINGARCH approach delivers much better results regarding goodness-of-fit and prediction when compared to the ordinary INGARCH model.
@@
We follow the arguments of Liu (2012) and show explicitly the steps required for the proof based on model (3); further details can be proved along the lines of Liu (2012) and are omitted.
By working in the same manner as in Liu (2012, pp.106) , we can prove that $\{\boldsymbol{\xi}_{t}\}_{t\in\mathbb{Z}}$ is a weak Feller chain and bounded in probability. Therefore, $\{\boldsymbol{\xi}_{t}\}_{t\in\mathbb{Z}}$ has at least one stationary distribution.
Let $w$ be a continuous function with compact support on $(0,\infty)^{2}$ and assume, without loss of generality, that it is bounded by 1, i.e. $|w|<1$. Consider $k=1$, ${\bf x}=(x_{1},x_{2}),{\bf z}=(z_{1},z_{2})\in(0,\infty)^{2}$, and $\epsilon>0$ arbitrary. Denote by $f^{mp}_{\boldsymbol{\xi};\mathcal{D}}(\cdot)$ the probability function of a mixed Poisson distribution with mean $\lambda$, variance $\lambda+\lambda^{2}/\phi$, and mixing distribution $\mathcal{D}$, where $\boldsymbol{\xi}=(\lambda,\phi)^{\top}$. It follows that
@@
$\sum_{y=0}^{\infty}|w(\boldsymbol{\tau}+{\bf B}(y,y)^{\top}+{\bf A}\boldsymbol{\bf x})f^{mp}_{\bf x;\mathcal{D}}(y)-w(\boldsymbol{\tau}+{\bf B}(y,y)^{\top}+{\bf A}\boldsymbol{\bf z})f^{mp}_{\bf z;\mathcal{D}}(y)|\leq$
$\sum_{y=0}^{\infty}f^{mp}_{\bf x;\mathcal{D}}(y)|w(\boldsymbol{\tau}+{\bf B}(y,y)^{\top}+{\bf A}\boldsymbol{\bf x})-w(\boldsymbol{\tau}+{\bf B}(y,y)^{\top}+{\bf A}\boldsymbol{\bf z})|+$ (A-1)
$\sum_{y=0}^{\infty}|w(\boldsymbol{\tau}+{\bf B}(y,y)^{\top}+{\bf A}\boldsymbol{\bf z})|\,|f^{mp}_{\bf x;\mathcal{D}}(y)-f^{mp}_{\bf z;\mathcal{D}}(y)|.$ (A-2)
We now find an upper bound for the term (A-2). Denote by $f^{pois}_{\lambda}(\cdot)$ the probability function of a Poisson distribution with mean $\lambda$. From the mixed Poisson stochastic representation, we have that $f^{mp}_{\bf x;\mathcal{D}}(y)=E(f^{pois}_{x_{1}Z_{1}}(y))$, where $Z_{1}$ is the associated latent random variables with distribution depending on $x_{2}$. Similar representation holds for $f^{mp}_{\bf z;\mathcal{D}}(y)$ in terms of an associated latent factor $Z_{2}$ with distribution depending on $z_{2}$. Using this and recalling the fact that $|\omega|<1$, we obtain that
$\sum_{y=0}^{\infty}|w(\boldsymbol{\tau}+{\bf B}(y,y)^{\top}+{\bf A}\boldsymbol{\bf z})|\,|f^{mp}_{\bf x;\mathcal{D}}(y)-f^{mp}_{\bf z;\mathcal{D}}(y)|\leq\sum_{y=0}^{\infty}|f^{mp}_{\bf x}(y)-f^{mp}_{\bf z}(y)|=$
$\sum_{y=0}^{\infty}|E(f^{pois}_{x_{1}Z_{1}}(y))-E(f^{pois}_{z_{1}Z_{2}}(y))|\leq E\left(\sum_{y=0}^{\infty}|f^{pois}_{x_{1}Z_{1}}(y)-f^{pois}_{z_{1}Z_{2}}(y)|\right).$ (A-3)
By using inequality (A.1) from Wang et al. (2014) (see also Liu (2012)), we have that $\sum_{y=0}^{\infty}|f^{pois}_{x_{1}Z_{1}}(y)-f^{pois}_{z_{1}Z_{2}}(y)|\leq 2\left(1-\exp\{-|x_{1}Z_{1}-z_{1}Z_{2}|\}\right)$. Hence, (A-1) is bounded above by
$E\left(\sum_{y=0}^{\infty}|f^{pois}_{x_{1}Z_{1}}(y)-f^{pois}_{z_{1}Z_{2}}(y)|\right)\leq 2E\left(1-e^{-|x_{1}Z_{1}-z_{1}Z_{2}|}\right)\leq$
$2E\left(1-e^{-|x_{1}-z_{1}|(Z_{1}+Z_{2})}\right)\leq 2E\left(|x_{1}-z_{1}|(Z_{1}+Z_{2})\right)=2|x_{1}-z_{1}|\leq 2\|{\bf x}-{\bf z}\|_{p},$